OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-4,2) and (1,-1).

OpenStudy (anonymous):

would it be \[y-2=\frac{ -3 }{ 5 }(x-4)\]

OpenStudy (anonymous):

 that is the answer i got

OpenStudy (anonymous):

@FibonacciChick666 @Nnesha @mathmath333 @ozzy22 @math92130

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

@gorv

OpenStudy (anonymous):

@JamesWolf

jimthompson5910 (jim_thompson5910):

you're very close

jimthompson5910 (jim_thompson5910):

but you made an error

jimthompson5910 (jim_thompson5910):

oh and I just caught the "slope intercept" so you have to convert over to y = mx+b form

OpenStudy (gorv):

x coordinate is -4

OpenStudy (gorv):

so we will right it as (x-(-4)) not the x-4

OpenStudy (gorv):

now we need to write in slope intercept form

OpenStudy (gorv):

u can conver your equation in that form

OpenStudy (anonymous):

ok

OpenStudy (gorv):

distribute -3/5 first insidethebracket

OpenStudy (anonymous):

ok um ill be right back

OpenStudy (gorv):

can u do that

OpenStudy (anonymous):

ok im back sorry about that

OpenStudy (gorv):

lol..u got what i told ??

OpenStudy (anonymous):

no i dont get do i have to put the fraction into brackets like these { or these ]

OpenStudy (gorv):

2(x+2)=2*x+2*2=2x+4 this how we distribute just multiply the number outside the bracket with element inside the bracket

OpenStudy (anonymous):

so the anwser is 6

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how are you getting 6?

OpenStudy (anonymous):

well um isnt it 2*1+4=6

OpenStudy (anonymous):

1=x

jimthompson5910 (jim_thompson5910):

why are you plugging in x = 1? and the expression 2(x+2) was just an example on how to distribute

OpenStudy (anonymous):

oh lol sorry

OpenStudy (anonymous):

so what would the anwser be @gorv

OpenStudy (gorv):

\[y-2=\frac{ -3 }{ 5 }(x+4)\]

OpenStudy (gorv):

\[y-2=\frac{ -3 }{ 5 }*x+4*\frac{ (-3) }{ 5 }\]

OpenStudy (gorv):

now add 2 on both side

OpenStudy (anonymous):

oh i made it - while it was supposed to be +

OpenStudy (gorv):

just add 2 on both side ...

OpenStudy (anonymous):

so y=3/s*x+4*1.4

OpenStudy (anonymous):

am i right or wrong

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

-3/5 = -0.6 you can't say 4*(-0.6) + 2 = 4*1.4 that's incorrect

jimthompson5910 (jim_thompson5910):

sure -0.6 + 2 = 1.4 is true, but that 4 up front changes things

OpenStudy (anonymous):

so the anwesr would be

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

You can't expect us to just give you the answer. Try it again.

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

hopefully you see how gorv got \[\Large y-2=\frac{ -3 }{ 5 }*x+4*\frac{ (-3) }{ 5 }\] right? or no?

OpenStudy (anonymous):

yea but i did it and thats how i got

jimthompson5910 (jim_thompson5910):

the next thing to do is to add 2 to both sides

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

y=-3/5*x+4 and how do i subtract -3/5 from 2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

first off, what is 4 times -3/5 equal to

OpenStudy (anonymous):

-2.4

jimthompson5910 (jim_thompson5910):

ok we can keep it as a decimal for now

jimthompson5910 (jim_thompson5910):

add 2 to that to get what number?

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

@omg1help y=-3/5*x+4 The 4 should be -(2/5)

jimthompson5910 (jim_thompson5910):

-2.4 + 2 is not 0.4

OpenStudy (anonymous):

sorry meant -0.4

jimthompson5910 (jim_thompson5910):

so we have (-3/5)x + (-0.4) or (-3/5)x - 0.4

jimthompson5910 (jim_thompson5910):

convert 0.4 to a fraction to get what?

OpenStudy (anonymous):

4/10

jimthompson5910 (jim_thompson5910):

reduce that

OpenStudy (anonymous):

2/5

jimthompson5910 (jim_thompson5910):

so the answer is \[\Large y = -\frac{3}{5}x - \frac{2}{5}\]

OpenStudy (anonymous):

really that was so easy i cant believe i didnt get it

jimthompson5910 (jim_thompson5910):

because that is equivalent to \[\Large y-2=-\frac{ 3 }{ 5 }*(x+4)\]